# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v4

@@ 6.2 Synthetic experiments on the selected railway line
    $Line1$, the double-track line with dense traffic. We use cyclic $3$-hour timetable with $10$ trains each hour and each direction, i.e. we have $59$ trains.
@@
    $Line2$, similar to $line1$, with the additional closure of one of the tracks between ZZ and RCB. We use a cyclic $2$-hour timetable with $10$ trains each hour and each direction, i.e. we have $40$ trains.
@@
    $Line3$, the whole line is considered as a single-track, with a timetable of $3$ hours traffic and $21$ trains.
For $line1$ and $line3$, the original timetable is feasible, while for $line2$ it is not feasible due to the closure. For each of the lines, we compute $12$ instances, each with different initial delays of train subsets. Instance 00 yields no initial delays. Thereafter, we use the parameter value $d_{\max}=40\ \text{min}$, in all cases in this and subsequent subsection. Assuming that each train interacts with $N(d_{\max})=12$ trains for $line1$ and $line2$, and with $N(d_{\max})=6$ trains for $line3$, Eqs. (17) - (19), (22) allows us to quickly estimate the sizes of the problem. These sizes are compared with the computed (comp.) actual number of variables and constraints (mean over instances) in Tab. 1. Here, $line2$ is the most complex in terms of constraints, which is reflected later by the computational difficulties of some of its instances. Furthermore, $line2$ is partially double-track and partially single-track, hence the actual number of constraints is bounded from above by the estimate for the single-track line and from below by the estimate for the double-track line. $Line1$ includes double-, whereas $line3$ single-track segments only. There actual number of constraints and variables appeared to be close to the estimate.
l #$\#\mathcal{S}$$\#\mathcal{J}$ int vars # precedence vars # constraints
comp. mean Eq. (17) act. mean
Eq. (18)/Eq. (20)
doub./sing.
comp. mean
Eq. (22)
doub./sing.
1 3 59 118 118 1538 1415 / - 7889 8731 / -
2 5 40 142 133 1394 1600/3200 8491 6867/13067
3 5 21 79 70 592 - / 840 3057 - / 3499
Table 1: Synthetic experiments. The first column is the ordinal number of the line. For the estimated and actual number of variables and constraints, we use $\alpha=2/3$ in Eqs. (17) - (19), (22) (doub. - double-track, sing. - single-track, comp. mean - computed actual number of variables or constraints, mean over instances).
@@
The computational results are presented in Fig. 3 a) b) c) for $line1$, $line2$, and $line3$, respectively. Each figure shows objective $\times$ $d_{\text{max}}$ (top part), computation time (comp. time, middle part), and QPU access time (bottom part) for CQM hybrid solver with t_min equal $5$ and $20$s. CPLEX results are also displayed.
@@
For scenario $line1$ (Fig. 3 a) ), the CQM solver obtained optimal solutions for all cases (except for instances $6$ and $11$). Similarly, for $line3$ (Fig. 3 c) ), only small deviations from the optimum can be observed for a number of instances. All computational times for $line1$ and $line3$ were significantly below $0.4$min, although CPLEX performed faster. Comparing CQM with t_min$=5s$ and t_min$=20s$, we have almost the same objectives but t_min$=20s$ requires almost $4$ times longer computational time.
@@
For $line2$ (Fig. 3 b) ), all CQM solutions were suboptimal, although feasible. More remarkably, for $9$ instances at t_min$=5s$ and $3$ instances at t_min$=20s$ (out of $12$ instances) CQM was faster than CPLEX, in several cases up to 5×5$\times$5 × faster. This difference in time is clear while setting the minimal processing time of the CQM solver t_min$=5s$ (see Section 5). In principle, $line1$ (Fig. 3 a) ) and $line3$ (Fig. 3 c) ) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster. We can conclude that the utility of the CQM solver can be expected for the more complex instances of railway scheduling problems.
@@
Figure 3: Synthetic experiments $Line1$, $Line2$, $Line3$ comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM. All the displayed solutions are feasible. Total computational time (middle panel) and QPU times (lower panel) were provided by the D-Wave output. Panel a) - $line1$, results of both solvers were similar in terms of the objective, but the CPLEX computation was faster. Panel b) - $line2$ namely the double-track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and for the quantum solver (in terms of objective). On this example, there are instances, where the current CQM D-Wave hybrid solver outperforms CPLEX in terms of computational time. Increasing the value of the t_min parameter in most cases improves the quality of the solution at the cost of computational time. Panel c) - $line3$ all the displayed solutions are feasible. Results of both solvers were similar in terms of the objective, but CPLEX computation was faster.
@@ 6.3 Real-life experiments on the considered rail network
In this subsection, we will test the CQM solver on more realistic scenarios of the railway traffic on the presented network. We will also focus on the track closure situation, turning a double-track line segment into a single-track one, under dense traffic, to elaborate our findings from $line2$ in Section 6.2. Our computations address various use cases of train delays outside the network and within the network as well as the partial track closure. We constructed $9$ networks with increased levels of difficulty ranging from small delays to disruptions, i.e. it is expected more trains are involved and/or disturbances are spread more over the network. These networks are tabulated in Table  2. Their details are as follows:
@@
    Networks with number $4$ or higher concern also disturbances due to closures.
@@
      Networks $4$ and $5$ concerns rerouting trains from double-track line KTC-CB to the single-track line with higher passing times, see Fig. 2 (trains have no initial delays in network $3$ and some initial delays in network $4$).
@@
      Network $6$ concerns multiple closures, i.e. change of multiple line KZ-KO(STM) and double-track line KO-KL to single-track lines. (See Fig. 2; the reason can be e.g. upcoming reconstruction works on this part of network.)
@@
      Networks $7$$9$ concerns closures of both $5$ and $6$, but each with different initial delays of trains. The intention of the last $3$ networks is to create a really challenging rescheduling problem.
@@ A. Disturbances from outside the network
2 10 arr. Ty KS 94766 KS 40518 IC 41004 KS 44862 IC 4120 Delayed 5 trains on the arrival to Ty, random delays of several minutes
@@ B. Disturbances originated within the network
3 10 whole network 14 trains KS + IC Random delays of 14 trains delays up to 35 minutes
@@ C. Disturbances within the network and closures
8 11 whole network all trains Closures as in 7 + random delays of 13 trains up to 30 minutes
9 11 whole network all trains Closures as in 7 + random delays of 15 trains up to 30 minutes
Table 2: Real-life experiments. Networks of particular dispatching problems. In networks 00 - $5$ most of analyzed network is double-tracked, from network $6$ upward the number of single-track lines in the network increases at the expense of the number of double and multiple-track lines. For each network we considered a two-hour time with an afternoon rush-hour timetable from the year $2021$, with #$\#J=27$ trains. Beside the #$\#S$ stations there are $2$ depots in each network.
@@
The level of difficulty increases with the number of variables and constraints. Observe that CQM always returns a feasible solution. For the networks $0-3$ with initial delays only, CQM generated optimal solutions, while in the remaining networks, the objective value was somewhat higher than the actual optima obtained with CPLEX. As for computational times, for CPLEX they grow significantly - from $0.072$ to almost $8s$ (network $9$). Meanwhile, those of CQM remain almost constant around $5.1s$. QPU times remain not greater than $0.032s$. Remarkably, for the hardest instances: networks $7$ and $9$, the CQM hybrid solver even outperforms CPLEX with respect to computational time. It does not find the actual optimum, just a feasible solution close to it. For example, in network $9$, the computed CQM solution is around 30% worse than the optimal one. In conclusion, we expect benefits from the application of the hybrid solver on medium-scale railway networks with multiple closures, as in networks 7-9. This finding follows our observation on synthetic data in Fig. 3 b), where the benefits appeared for more complex instances, too.
To evaluate the extent of difficulty of the problems analyzed in Tab. 3, Tab. 4 presents mean (over stations and realizations) delays, resulting from CQM solutions, on selected stations (i.e. stations in the central part of the network in Fig. 2). Meaningful delays are introduced starting from network $4$ onward, i.e. for networks with closures in a limited number of stations, see also Tab. 2. From network $5$ onward, delays become more uniformly distributed among the stations when compared with results in Tab. 2. However, for networks $7$ $9$ the mean delays are the highest and almost uniformly distributed among all stations, showing that the spread of disruption impacts the whole network. These have appeared to be more challenging network disruptions.
Fig. 4 shows a more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (left, red) vs CQM (right, blue). For network $4$, bigger delays occur mainly at stations KO(STM), KO and CB for both CPLEX and CQM solutions, albeit for the former are somewhat smaller. This reflects the initial disturbances, which indeed occurred mainly between KO and CB (see Tab. 2), and thus the spread is limited to these 3 stations. For network $7$, delays are more spread through the network due to disruption and initial disturbances. To conclude, we can expect an advantage from a hybrid (quantum-classical) approach, particularly in case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network. To facilitate further evaluation of the solutions, train diagrams are presented in Appendix A.
@@
Table 4: Real-life experiments. Mean (over trains and realizations) secondary delay resulted from CQM solutions (Tab. 3) for all networks in Tab. 2. Non-zero values for network 00 come from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes, delay but makes it up and leaves the network on time - these are typical issues for the railway traffic on analyzed network.
Figure 4: Real-life experiments. Statistics of CPLEX (left, read) and CQM (right, blue) solutions (Tab. 3) for selected networks in Tab. 2. The statistics were calculated over trains in the classical cases, and over trains and realizations in the CQM cases. Vertical lines are ranges, while horizontal lines are $0.25$ and $0.75$ percentiles.
The right choice of the t_min parameter can improve results meaningfully. Let us analyze in more detail the role of the t_min solver parameter for selected problems. In particular, for networks $7$ and $9$ therein, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Fig. 5 a) and b). The objective value is roughly log-linear in computational time. For small t_min parameter values the computational time is shorter, whereas for large t_min it grows exponentially. Overall, the improvement in the objective is overwhelmed by the increase of computational time. This is important from the point of view of the algorithm layout, as it may be tied to a power law scaling. Such behavior is plausible in the case of QA methods (Soriani et al., 2022; Albash et al., 2017; Domino et al., 2022a).
Figure 5: Real-life experiments. The sweep over t_min parameter for network $7$ a) and network $9$ b), see Tab. 2. All solutions were feasible. For higher t_min we expect smaller objective values but higher total computational time. For each parameter value, there were $5$ realizations of the experiment were performed. Interestingly, a linear (negatively sloped) relation between the objective and the logarithm of ${\tt t\_min}$ can be observed, suggesting power law scaling.
@@ 7 Conclusions
We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks. In particular, we have encoded rescheduling problems in a linear integer program and solved these using a state-of-the-art solver (CPLEX), as well as the QA based CQM hybrid (quantum-classical) solver. The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready to tackle challenging railway rescheduling problems. While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were actually better. This supports the expectation that future quantum devices will be efficient in solving large-scale problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that is beyond the scope of current exact models and heuristics.
While the quantum-based solvers possibly return suboptimal solutions, these are still feasible and often better than the solutions obtained manually or based on smaller-scale models. In each computation, the CQM hybrid solver always has taken some advantage of the QPU. Hence, such a little help from the QPU must have boosted the classical heuristics in the hybrid solver, even if quantum annealer devices are subjects of many ongoing debates.
Several future research directions can be determined. The first concerns the ongoing worldwide works on the development of fault-tolerant quantum computing devices (DiVincenzo, 2000) with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular). The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of quantum device to our advantage e.g. by modeling some stochastic behavior on railways. The third concerns the evolution and application of quantum-inspired techniques such as tensor networks technique, and simulated bifurcation (Goto, 2019). With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.

### v5

@@ 6.2 Synthetic experiments on the selected railway line
    $line1$, the double-track line with dense traffic (As illustrated in Fig 2, we have $3$ decision stations: KO, CB and GLC). We use 3 periods of an hourly repeating timetable with $10$ trains each hour and in each direction. Here, and also in the following lines we did not consider one of the trains for practical reasons, thus we have $59$ trains in a three-hour time horizon.
@@
    $line2$, similar to $line1$, with the additional closure of one of the tracks between ZZ and RCB, hence, these stations are added to the set of decision stations, and single-track traffic is assumed between them. We have $39$ trains on a two-hour time horizon.
@@
    $line3$, the whole line is considered a single track. As in the "$line2$" case, we add two extra decision stations, ZZ and RCB where M-P and M-O are possible. We have $21$ trains in a three-hour time horizon.
For $line1$ and $line3$, the original timetable would be feasible without the initial delays, while for $line2$ it is not feasible due to the closure. The instances are summarized in Tab 4.
line #$\#J$ n.o. hours decision stations track layout
1 $59$ $3$ KO, CB, GLC double
2 $36$ $2$ KO, CB, RCB, ZZ, GLC
single between RCB and ZZ
then double
3 $21$ $3$ KO, CB, RCB, ZZ, GLC single
Table 4: Features of synthetic instances.
For each of the lines, we compute $12$ instances, each with different initial delays of train subsets. The initial delays were assigned manually, a-priory for each instance, choosing plausible values based on everyday practice. Instance 00 has no initial delays. Thereafter, we use the parameter value $d_{\max}=40\ \text{min}$, in all cases in this and subsequent subsection. We chose its particular values by experimentation.
The upper limits found in Section 4.2.3 are compared to the actual number of variables and constraints (mean over instances) in Table 5. The purpose of the table is to connect the theoretical discussion on scaling in Section 4.2.3 with the actual problem sizes. For example, $line2$ is the most complex in terms of constraints, which is reflected later by the computational difficulties of some of its instances. (Due to closure, $line2$ is partially double-track and partially single-track with heavy traffic.). The table serves also as the sanity check of our model.
$line$$\#\mathcal{S}$$\#\mathcal{J}$ int vars # precedence vars # constraints
actual mean
Equation (20)
upper lim
actual mean
Equation (21)
upper lim
acctual mean
Equation (25)
upper lim
1 3 59 118 180 1538 5221 7889 24721
2 5 39 142 200 1394 3802 8491 18642
3 5 21 79 110 592 1102 3057 5502
Table 5: Estimated and actual problem sizes on $line1$ $line2$ and $line3$ scenarios. For each scenario, the mean over instances was computed.
@@ 3 $21$ $3$ KO, CB, RCB, ZZ, GLC single
The computational results are presented in Figure 3 a) b) c) for $line1$, $line2$, and $line3$, respectively. Each figure shows the objective (top part), computation time (comp. time, middle part), and QPU access time (bottom part) for CQM hybrid solver with t_min equal $5$ and $20$s. CPLEX results are also displayed. These latter include the optima as well as the heuristic results with computational time limited to at most 5 seconds.
@@
For the scenario $line1$ (Figure 3 a) ), the CQM solver found the optimal solutions in all of the cases except for instances $6$ and $11$. Similarly, for $line3$ (Figure 3 c) ), only small deviations from the optimum can be observed for a number of instances. All computational times for $line1$ and $line3$ were significantly below $30$seconds, although CPLEX was faster, and found the actual optimum before the possible 5 seconds limit. Comparing CQM with t_min$=5s$ and t_min$=20s$, we have almost the same objective values, however, t_min$=20s$ requires almost $4$ times longer computational time.
@@
For $line2$ (Figure 3 b) ), all CQM solutions were suboptimal, although feasible. More remarkably, for $9$ instances at t_min$=5s$ and $3$ instances at t_min$=20s$ (out of $12$ instances) CQM was faster than exact CPLEX, in several cases up to $5$ times faster. CPLEX could not solve these instances to optimality when it was limited to 5 seconds; it returned close-to-optimal feasible solutions though. This difference in time is clear while setting the minimal processing time of the CQM solver t_min$=5s$ (see Section 5). In principle, $line1$ (Figure 3 a) ) and $line3$ (Figure 3 c) ) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster. We can conclude that the utility of the CQM solver can be expected for the more complex instances of railway scheduling problems, as the present instances for $line2$.
The detailed analysis of the implications of suboptimality in CQM solutions for $line2$ from the railway point of view shows that there are two types of consequences: the increase of the passing time between subsequent nodes, and the change in certain trains’ order. Concerning the first issue, the realization of such suboptimal solutions can result in slowing down trains, which can be a motivation for further research, such as in the work of [14]. Solutions affected by the second issue are suboptimal but still feasible and can be opted for by the human dispatcher because they have benefits for certain non-quantified reasons. Particular examples of real-life problem instances are presented in Appendix A.
The hybrid quantum-classical solver CQM does not outperform purely classical heuristics in most problems at the present state of the art, and railway dispatching is not yet an exception. Our results demonstrate, however, that it already has the potential to handle difficult railway problems of moderate size. This implies that CQM, as a generic out-of-the-box solver, could already be a viable alternative to the currently existing solvers. Note that suboptimality can be accepted in many cases as it will be justified by the analysis of our real-world examples in Section 6.3. Further improvement in the runtime and the solution quality can possibly be achieved by developing custom classical-quantum heuristics tailored for the given problem, providing also information on when the quantum solvers can be the most efficient. Meanwhile, as there is a significant R&D effort put into the development of quantum hardware to get beyond the NISQ era, custom algorithms, and even generic hybrid solvers can become more competitive in the future. This may render quantum approaches important or even dominant for more complex disruptions and on larger scale.
@@
Figure 3: Synthetic experiments $line1$, $line2$, $line3$. Comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM. All the displayed solutions are feasible. Total computational time (middle panel) and QPU times (lower panel) were provided by the D-Wave output. Panel a) - $line1$, results of both solvers were similar in terms of the objective, but the CPLEX computation was faster. Panel b) - $line2$ namely the double-track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and the quantum solver (in terms of objective). In this example, there are instances where the current CQM D-Wave hybrid solver provides feasible, though not optimal solutions of similar quality to those that CPLEX provides when its computational time is limited to be similar to that of CQM. When CPLEX is required to solve the problem to optimality, it can be slower than CQM. Increasing the value of the t_min parameter does, in most cases, improve the quality of the solution at the cost of computational time. Panel c) - $line3$ all the displayed solutions are feasible. The results of both solvers were similar in terms of the objective, but CPLEX computation was faster.
@@ 6.3 Real-life experiments on the considered rail network
In this subsection, we will test the CQM solver on more realistic scenarios of the railway traffic on the presented network. We will also focus on the track closure situation, turning a double-track line segment into a single-track one, under dense traffic, to elaborate our findings from $line2$ in Section 6.2. Our computations address various use cases of train delays outside the network and within the network as well as the track closures. We constructed $9$ networks with increased levels of difficulty ranging from small delays to disruptions, i.e. it is expected more trains are involved and/or disturbances are spread more over the network. Delays were assumed a priority fixed for each network manually according to everyday practice. These networks are tabulated in Table  6. Their details are as follows:
@@
    Networks with number $4$ or higher also include disturbances due to closures.
@@
      Networks $4$ and $5$ concern rerouting trains from the double-track line KTC-CB to a single-track line with higher passing times, see Figure 2 (trains have no initial delays in network $3$ and some initial delays in network $4$).
@@
      Network $6$ concerns multiple closures, i.e. change of multiple line KZ-KO(STM) and double-track line KO-KL to single-track lines. (See Figure 2; the reason can be e.g. upcoming reconstruction works on this part of the network.)
@@
      Networks $7$ $9$ concern closures of both $5$ and $6$, but each with different initial delays of trains. Our intention with the last $3$ networks is to create a truly challenging rescheduling problem.
@@ A. Disturbances from outside the network
2 10 arr. Ty KS 94766 KS 40518 IC 41004 KS 44862 IC 4120 Delayed 5 trains on the arrival to Ty, arbitrary delays of several minutes
@@ B. Disturbances originated within the network
3 10 whole network 14 trains KS + IC Arbitrary delays of 14 trains delays up to 35 minutes
@@ C. Disturbances within the network and closures
8 11 whole network all trains Closures as in 7 + arbitrary delays of 13 trains up to 30 minutes
9 11 whole network all trains Closures as in 7 + arbitrary delays of 15 trains up to 30 minutes
Table 6: Real-life experiments. Networks of particular dispatching problems. In networks 00 - $5$ most of the analyzed network is double-tracked, from network $6$ upward the number of single-track lines in the network increases at the expense of double and multiple-track lines. In each experiment instance, initial delays were chosen manually, as arbitrary values plausible in practice. For each network, we considered a two-hour time with an afternoon rush-hour timetable from the year $2021$, with #$\#J=27$ trains. Besides the #$\#S$ stations there are $2$ depots in each network.
@@
The level of difficulty increases with the number of variables and constraints. Observe that CQM always returns a feasible solution. For networks $0-3$ with initial delays only, CQM generated optimal solutions, while in the remaining networks, the objective value was somewhat higher than the actual optima obtained with CPLEX. As for computational times, for CPLEX they grow significantly - from $0.072$ to almost $8s$ (network $9$). Meanwhile, those of CQM remain almost constant at around $5.1s$. QPU times remain not greater than $0.032s$. Remarkably, for the hardest instances: networks $7$ and $9$, the CQM hybrid solver provides feasible solutions faster than exact CPLEX. It does not find the actual optimum though, just a feasible solution close to it. Certainly, there can be other (meta)heuristics that provide solutions of similar quality in an even shorter time, for instance, CPLEX itself when running for a limited time, as we have found in the case of the synthetic examples in Section 6.2. (This experience is also iniline with cases of other similar scheduling problems, see the work of [33] for another example.) Nevertheless, the experience with these examples suggests that CQM provides useful results in reasonable time, so it is at least one of the possible heuristics worth considering. We can expect benefits from the application of the hybrid solver on medium-scale railway networks with multiple closures, as in networks 7-9. This finding follows our observation on synthetic data in Figure 3 b), where the benefits appeared for more complex instances, too.
To evaluate the extent of difficulty of the problems analyzed in Table 7, Table 8 presents the values of the delays averaged over trains and realizations, resulting from CQM solutions, for selected stations (i.e. stations in the central part of the network in Figure 2). Delays with a broader impact are introduced starting from network $4$ onward, i.e. for networks with closures in a limited number of stations, see also Table 6. From network $5$ onward, delays become more uniformly distributed among the stations when compared with results in Table 6. However, for networks $7$ $9$ the mean delays are the highest and almost uniformly distributed among all stations, showing that the spread of disruption impacts the whole network. These appeared to be more challenging network disruptions.
Figure 4 shows a more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (left, red) with exact CQM (right, blue). For network $4$, bigger delays occur mainly at stations KO(STM), KO, and CB for both CPLEX and CQM solutions. This reflects the initial disturbances, which indeed occurred mainly between KO and CB (see Table 6), and thus the spread is limited to these 3 stations. For network $7$, delays are more spread through the network due to disruption and initial disturbances. To conclude, we can expect an advantage from a hybrid (quantum-classical) approach, particularly in the case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network. To facilitate further evaluation of the solutions, train diagrams are presented in Appendix A.
@@
Table 8: Real-life experiments. Average (over trains and realizations) secondary delay resulted from CQM solutions (Table 7) for all networks in Table 6. Non-zero values for network 00 come from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes, delay but makes it up and leaves the network on time - these are typical issues for the railway traffic on the analyzed part of the network.
Figure 4: Real-life experiments. Statistics of CPLEX (left, read) and CQM (right, blue) solutions (Table 7) for selected networks in Table 6. The statistics were calculated over trains in the classical cases, and over trains and realizations in the CQM cases. Vertical lines are ranges, while horizontal lines are $0.25$ and $0.75$ percentiles.
The right choice of the t_min parameter can improve results meaningfully. Let us analyze in more detail the role of the t_min solver parameter for selected problems. In particular, for networks $7$ and $9$, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Figure 5 a) and b). The objective value is roughly log-linear in computational time. For small t_min parameter values the computational time is shorter, whereas for large t_min it grows exponentially. Overall, the improvement in the objective is overwhelmed by the increase in computational time. This is important from the point of view of the algorithm layout, as it may be tied to a power law scaling. Such behavior is plausible in the case of QA methods [34, 1, 16].
Figure 5: Real-life experiments. Dependence on the t_min parameter for network $7$ a) and network $9$ b), see Table 6. All solutions were feasible. For higher t_min we expect smaller objective values but higher total computational time. For each parameter value, there were $5$ realizations of the experiment performed. Interestingly, a linear (negatively sloped) relation between the objective and the logarithm of ${\tt t\_min}$ can be observed, suggesting power law scaling.
@@ 7 Conclusions
We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks. In particular, we have encoded rescheduling problems in a linear integer program and solved these using a state-of-the-art solver (CPLEX), as well as the QA based CQM hybrid (quantum-classical) solver. The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready to tackle challenging railway rescheduling problems. While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were competitive. This supports the expectation that future quantum devices will be efficient in solving large-scale problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that is beyond the scope of current exact models and heuristics.
While the quantum-based solvers possibly return suboptimal solutions, these are still feasible and often better than the solutions obtained manually or based on smaller-scale models. In each computation, the CQM hybrid solver always has actually used the QPU. Hence, such a little help from the QPU may have boosted the classical heuristics in the hybrid solver, even if quantum annealer devices are subjects of many ongoing debates.
Several future research directions can be determined. The first concerns the ongoing worldwide works on the development of fault-tolerant quantum computing devices [13] with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular). The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of the quantum device to our advantage e.g. by modeling some stochastic behavior on railways. The third concerns the evolution and application of quantum-inspired techniques such as tensor networks technique, and simulated bifurcation [20]. With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.